### v1

# Optimising quantum tomography via shadow inversion

Abstract
In quantum information theory, the accurate estimation of observables is pivotal for quantum information processing, playing a crucial role in compute and communication protocols. This work introduces a novel technique for estimating such objects, leveraging an underutilised resource in the inversion map of classical shadows that greatly refines the estimation cost of target observables without incurring any additional overhead. A generalised framework for computing and optimising additional degrees of freedom in the homogeneous space of the shadow inversion is given that may be adapted to a variety of near-term problems. In the special case of local measurement strategies we show feasible optimisation leading to an exponential separation in sample complexity versus the standard approach and in an exceptional case we give non-trivial examples of optimised post-processing for local measurements, achieving the same efficiency as the global Cliffords shadows.
Introduction
Quantum state tomography with classical shadows [1, 2], initially motivated by the seminal work of Aaronson [3], represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems. This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications [4, 5, 6, 7], which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead[8, 9, 10, 11, 12, 13, 14, 15, 16]. At the core of it lies the concept of constructing these aforementioned shadows; efficiently representable compressed classical descriptions of quantum states. These shadows are generated through a process involving random measurements on copies of a state, followed by classical processing to reconstruct a succinct representation of the state, the ’shadow‘. These classical shadows, though not providing a complete description of the quantum state, contain enough information to estimate a wide range of properties with high accuracy when an appropriate POVM is used. An archetypal example is the set of measurements performed by random perturbations sampled from the Clifford group [17, 18] followed by a computational basis measurement
$\mathcal{C}_{N}=\{T\in SU(2^{N})|TP_{N}T^{\dagger}\in\mathcal{P}_{N}\}/U(1),$ (1)
for all $P_{N}$, the set of Pauli operators on $N$ qubits. Critically for shadow tomography, we have that the Clifford group constitutes a $2,3$-design [17] i.e., it can reproduce the statistical properties of the full set of unitaries up to the second and third moment. It is this that forms the basis upon which shadow tomography derives its impressive performance, implementing a map whose consequent inverse lies at the core of efficient estimation with classical shadows. If $U_{k}$ are elements of the $N$ qubit Clifford group and $\Pi_{\ell}$ are computational measurement projectors all acting on an input $N$ qubit quantum state then
$\mathcal{M}[\rho]=\frac{1}{|\mathcal{C}_{N}|}\sum_{k=1}^{|\mathcal{C}_{N}|}\sum_{\ell=1}^{2^{N}}T_{k\ell}\rho T_{k\ell}=\lambda\rho+(1-\lambda)\frac{\mathbbm{1}}{2^{N}},$ (2)
for $T_{k\ell}=U_{k}^{\dagger}\Pi_{\ell}U_{k}$. The titular inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$. Classical shadows great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum. This rapid convergence to a known configuration means the linear map of (2) may be readily inverted to recover the original state $\rho$. As shown in [2], the inverse map of (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
Figure 1: (a) Classical shadow tomography with optimised shadow inversion. After a collection of $M$ POVM outcomes have been acquired from a fixed quantum state, any estimation can be improved for a dynamic choice of target observable $A$ by computing an optimal ‘inverse’ operation for that observable and POVM. (b) Shadow norm for different strategies. The shadow norm cofactor when optimised for certain target projectors, using local Pauli measurements on single qubits. The shadow norm for such an observable on $N$ qubits is normally $(3/2)^{N}$ but when optimised over additional degrees of freedom can be shrunk to $(1.15)^{N}$. Though still exponential in the number of qubits, the optimised strategy is exponentially cheaper than standard classical shadow tomography with local measurements.
In these cases, the inversion map takes a particularly simple form, and the estimation of consumed resources (sampling complexity) critically depends on it. Nevertheless, what is often neglected in studies is the fact that such a map is not unique, i.e., as long as the POVM used is overcomplete (in the operator, Hilbert-Schmidt space), an additional set of free parameters appear in the inversion process which may be optimised to improve the post-processing performance of the estimation procedure. This is precisely the main task of this work – we will show the general inverse map contains a free homogeneous component $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ that may be further optimised. Our goal is to pursue such an optimisation by considering a hereto before missed resource in the tomography procedure that appears most vividly in classical shadows and the post-processing thereof. Though the natural choice of the inverse map (such as the one given above) is the inverse of the forward channel being applied, there is nothing that says it must be so. It is this inversion process, normally treated as a fixed mathematical procedure, that we aim to modify and improve upon. As long as the set of performed measurements is tomographically over-complete, the additional degrees of freedom may be exploited to reduce case-specific variance for a given observable. In doing so, we will show that an exponential improvement in estimations is possible compared to the standard choice of inverse map presented in the literature. With this novel approach to processing acquired quantum data, we open a new route to exploit areas where advantages may be gained without incurring an unacceptable resource cost.
Shadow map inversion
We begin by describing a generalised procedure, depicted in Figure 1(a) fulfilling the random measurement scheme with a fixed POVM $\{E_{k}\}_{k=1}^{n}$ [19, 20] on a $d$-dimensional system. The POVM elements sample randomly an unknown state $\rho$ with probability $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$. These POVM elements span a subspace $\mathcal{V}_{D}=\mathrm{Span}\{E_{k}\}_{k=1}^{n}$ in the Hilbert-Schmidt space (HS) of the dimension $|\mathcal{V}_{D}|=D$. Typically, $D=d^{2}$ (the entire HS space), but for the sake of generality, we shall keep $D\leq d^{2}$. Furthermore, we assume $D<n$, thus $E_{k}$ forms an over-complete basis in $\mathcal{V}_{D}$. The estimation task aims to approximate the mean values of an observable $A\in\mathcal{V}_{D}$, i.e., $\expectationvalue{A}=\mbox{Tr}\left[{A\rho}\right]$. Note that $\expectationvalue{A}=\mbox{Tr}\left[{A\bar{\rho}}\right]$, with $\bar{\rho}$ being the projection of a quantum state $\rho$ onto $\mathcal{V}_{D}$ which follows directly from $A\in\mathcal{V}_{D}$. For the case of the entire HS space, we have $\bar{\rho}=\rho$. Similarly, the measured probabilities trivially satisfy $p_{k}=\mbox{Tr}\left[{E_{k}\rho}\right]=\mbox{Tr}\left[{E_{k}\bar{\rho}}\right]$. The post-processing task assigns a (matrix) estimator $\hat{\eta}\in\{\dots,\hat{\eta}_{k},\dots\}$, i.e. each outcome $E_{k}$ the corresponding matrix $\hat{\eta}_{k}$, which represents a single-shot approximation of $\bar{\rho}$. They can be calculated via an inversion map $\hat{\eta_{k}}=\mathcal{M}^{-1}(E_{k})$ [2, 20] which we will show is not unique, i.e. $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ with the homogeneous component containing auxiliary parameters that may be used as a free resource in post-processing.
The overall classical shadow itself is obtained by averaging over the estimators obtained from $M$ random measurements: $M^{-1}\sum_{s}^{M}\hat{\eta}^{(s)}$. Assuming that all samples are i.i.d., the average value $\expectationvalue{\hat{\eta}}$ is immediately
$\expectationvalue{\hat{\eta}}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}\overset{!}{=}\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
This is the requirement on the estimator $\hat{\eta}$ such that the statistics of the shadow mimics, in expectation, that of the actual unknown state. Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, any solution to Eq. (3) is not unique. This multiplicity results in a set of free parameters, which becomes an extra resource that may be used in the post-processing component of an estimation to reduce the variance.
To achieve this, we now show a constructive method to recover families of equivalent classical shadows in terms of the POVM elements and their free parameters. We rely on an equivalent problem to (3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
$\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]\overset{!}{=}\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
It is instructive to express the operators as vectors in $\mathcal{V}_{D}$, i.e., for an operator $X\in\mathcal{V}_{D}$, we shall use the ket notation $|{X}\rangle\!\rangle$. We also identify an orthonormal basis $\beta_{D}=\{|{B_{j}}\rangle\!\rangle\}_{j=1}^{D}$ in $\mathcal{V}_{D}$. The expectation value from (4) can be expressed as
$\expectationvalue{\,a\,}=\sum_{k=1}^{n}\langle\!\langle{A^{\dagger}}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|{\rho}\rangle\!\rangle\overset{!}{=}\langle\!\langle{A^{\dagger}}|{\bar{\rho}}\rangle\!\rangle$ (5)
where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$. To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$
$\sum_{k=1}^{n}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|=\sum_{k=1}^{n}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|\equiv\mathbbm{1}_{D},$ (6)
which can be solved for $|{\eta_{k}}\rangle\!\rangle$ using $E_{k}^{\dagger}=E_{k}$. The last equation expressed in the basis $\beta_{D}$ reads
$\sum_{k=1}^{n}\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|{B_{t}}\rangle\!\rangle=\sum_{k=1}^{n}L_{sk}R_{tk}^{*}\overset{!}{=}\delta_{st},$ (7)
with $L_{sk}=\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle$ and $R_{tk}=\langle\!\langle{B_{t}}|{E_{k}}\rangle\!\rangle$. This implies the matrix equation $LR^{\dagger}=\mathbbm{1}_{D}$; thus, $L$ is a generalised inverse of $R^{\dagger}$. Since $R$ is a $D\times n$ matrix having $|{E_{k}}\rangle\!\rangle$ as columns, it is a rank-$D$ matrix (recall there are only $D$ independent $E_{k}$ matrices). We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero-matrix. Having this, a simple exercise shows $L=W(\Sigma^{-1}|H)U^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix. Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=W(\Sigma^{-1}|\hat{0})U^{\dagger}$ (particular) and $L^{(h)}=W(\hat{0}|H)U^{\dagger}$ (homogeneous). This results in a family of equivalent estimators
$\hat{\eta_{k}}=\sum_{s=1}^{D}L_{sk}\hat{B_{s}}=\sum_{s=1}^{D}\left(L_{sk}^{(p)}+L_{sk}^{(h)}\right)\hat{B_{s}}=\hat{\eta}_{k}^{(p)}+\hat{\eta}_{k}^{(h)},$ (8)
parameterised by a homogeneous solution $\hat{\eta}_{k}^{(h)}$. As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimisation in post-processing. We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$ where the homogeneous coefficients do not influence expectation values or any other linear combination of classical shadows
$\expectationvalue{\,a\,}=\sum_{k=1}^{n}p_{k}a_{k}=\sum_{k=1}^{n}p_{k}\left(a_{k}^{(p)}+a_{k}^{(h)}\right)=\sum_{k=1}^{n}p_{k}a_{k}^{(p)}.$ (9)
By substituting $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{AE_{k}}\right]$ into the previous equation we find
$\sum_{k}a_{k}^{(p)}E_{k}=A,$ (P)
$\sum_{k}a_{k}^{(h)}E_{k}=0.$ (H)
The set of solutions to the latter (homogeneous) equation defines an auxiliary optimisation space of dimension $n-D$ independent of the target observable (fixed by the choice of POVM). This is a free resource that may now be used in post-processing.
Figure 2: (a) Optimal free parameter for projectors on the Bloch sphere. The contour plot shows the minimal shadow norm (11) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$). Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map (13) found in [2] (b) Optimal free parameter for plane projectors. The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$). This enhanced norm is always lower or equal to the standard shadow norm |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\leq 5/4$ (see main text). In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis. The bound corresponds to the combinations which present optimal variance |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy. It is customary in classical shadow tomography to refer to variance as figure of merit [2]. In particular, we consider the state-independent upper bound on variance, defined as the shadow norm, which only depends on the decomposition of the particular observable:
$\mbox{Var}\left[{a}\right]\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{shadow}}.$ (10)
This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably. This is akin to the original upper bound on shadow norm and thence sampling complexity found in [2], where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{shadow}})$ ). Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator. By optimising over these coefficients, it is possible to identify the configuration which guarantees the minimal sample size for the estimation of the particular observable $A$:
$\|A\|^{2}_{\text{opt}}=\min_{a_{k}^{(h)}:H}\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right],$ (11)
The existence of a minimum is guaranteed through the positive semidefiniteness of the variance operator. We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in (11) yields exceptional gains for relatively minuscule cost.
Product observables and shadow optimisation
Let us now discuss the estimation of product observables $A^{(N)}=A_{1}\otimes A_{2}\ldots\otimes A_{N}$ via local POVMs, i.e., of the type $E_{k_{1}}\otimes\dots\otimes E_{k_{n}}$. This case is particularly interesting for near-term applications, such as the VQE algorithm [21, 22]. In such a scenario, it is easy to show that the shadow norm in Eq. (10) has a product form (see also [20]) and consequently Eq. (11) becomes
$\|A^{(N)}\|^{2}_{\text{opt}}=\prod_{j=1}^{N}\|A_{j}\|^{2}_{\text{opt}}$ (12)
The optimisation in Eq. (10) appears efficient, given the total number of free parameters scales linearly with the number of subsystems $N$. We see that even minimal deviation from the minimum in local shadow norms $\|A_{j}\|^{2}_{\text{opt}}$ may accumulate exponentially fast with $N$. This is illustrated in Figure 1(b) for the case of the canonical inversion map of Eq. (13). We now present examples of such exponential improvement for qubits for particular sets of local observables, specifically projectors, whereby by adjusting the available free parameters, we are able to reduce the expected bound on sample size based on assumptions for the observable itself.
Fidelity estimation of product states
We first consider the case of estimation of fidelity with product states. The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterised on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$. The measurement scheme is described by the POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$. As reference, we consider the bound $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{shadow}}\equiv(3/2)^{N}$ obtained using the estimators
$\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{3}\right).$ (13)
found in [2]. Now we show how to improve this bound by our techniques. The POVMs $E^{\pm}_{\zeta}$ span the entire HS space. We have $n=6$ and $D=4$, thus the equation (H) gives two free optimisation parameters for each qubit (details can be found in appendix A). In FIG.2(a), we present the minimal upper bound as a function of angles $(\theta,\phi)$, which is contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$. Since the overall shadow norm (12) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
Locally efficient fidelity estimation implies globally efficient
We now present a configuration in which the optimisation of the inversion map not only improves the standard bound, but we are actually able to achieve efficiency, i.e. $\|A\|^{2}_{\text{opt}}\equiv 1$. We consider the POVM constituted by the projectors of the X and Y Paulis, which spans the ”equator” on the Bloch sphere. The target observables are tensor products of a projector on a single qubit, parameterised by a single angle: $A=\outerproduct{\psi(\phi)}{\psi(\phi)}$ | with $|{\psi(\phi)}\rangle=(|{0}\rangle+e^{i\phi}|{1}\rangle)\sqrt{2}$. As before, we follow the ansatz given in [2] to get the particular solution for the inversion map and estimators
$\hat{\eta}^{\pm}_{\zeta}=4E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}.$ (14)
From here we obtain the reference bound for the shadow norm of $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\equiv(5/4)^{N}$. We reference to Appendix B.0.2 for more precise remarks on how we obtain this result.
Now we turn to our optimisation technique. We have the POVM with $n=4$ elements spanning a subspace of dimension $D=3$, therefore the generic estimators obtained from our procedure depend on a single free parameter $p$. The shadow norm of single-qubit projector depends on this free parameter and on the angle parameter $\phi$:
|$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}=\frac{1}{2}$ $(1+p_{i}^{2}+B(\phi,p_{r})+$ (15)
$\quad\sqrt{2}\sqrt{C(\phi,p_{r})})$
[…]
$\quad+2p_{r}(p_{r}-1)$
$C(\phi,p_{r})=(4$ $)\displaystyle p_{r}-1)\cos{\phi}+2p_{r}\cos{2\phi}$ ) (17)
$\quad+\cos{3\phi}+2+2p_{r}(p_{r}-1),$
[…]
In general, the dependency on $\phi$ remains explicit, changing the expected bound based on the projector to be estimated. In FIG.2(b) we present the optimal value of the shadow norm for each parameter $\phi\in[0,\pi]$. We note that for each value of $\phi$ the optimised norm is |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{opt}}\leq 5/4$. Moreover, it can be seen that for specific angles, the bound is exactly $\|E_{\zeta}^{\pm}\|^{2}_{\text{opt}}\equiv 1$., i.e., for the eigenprojectors of $X$ and $Y$ Paulis, corresponding to the POVM elements themselves. This implies efficient estimation: given the product form of the overall shadow norm (12), this scaling represents not only an exponential improvement compared to $(5/4)^{N}$, as for other projectors, but also an expected constant scaling of the sample size with the number of qubits. This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $2^{N}$ projections simultaneously for complementary observables. Previously this result has been found only via global measurement strategies relying on Clifford unitaries consisting of many joint operations, acting on all $N$ qubits at the same time; we have shown instead that by introducing the initial assumption over the space of observables, this could be achieved using only local single-qubit measurements.
Discussion
In the course of this letter we have observed how post-processing, an often overlooked aspect in estimation, represents a relevant resource to be examined in the accounting of a more efficient tomography. By introducing a generalised framework to obtain families of equivalent parameter-dependent estimators, we have found a clear relation between over-complete POVMs and their ability to reduce the variance in estimation of arbitrary observables. When considering local measurements on large composite quantum systems, even a small gain on each estimator represents an overall improvement in sample complexity, yielding an exponential separation between the naive sample variance and the optimised one as seen in Figure 1(b). In the case of example two, this optimised gain can achieve a variance that performs as well as global classical shadows for given observables, with modest additional post processing that remains efficient. We anticipate that by considering overcomplete POVMs containing joint measurements on multiple qubits, further advantage may be found at increased (but still tractable) optimisation difficulty, either in sample complexity or the set of observables. Finally, interleaving optimised multiqubit measurements may yield sample improvements that approach regular classical shadows within some limit. Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimisation will be the topic of future work.
Acknowledgements
We thank Richard Küng for his insight and excellent discussions on the subject matter. This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/F7115] (BeyondC). For the purpose of open access, the author(s) has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Appendix A Parameter dependence for estimators
The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$, admits a homogeneous combination $\sum_{i=\pm}\sum_{\zeta}h_{i,\zeta}E^{i}_{\zeta}=0$ (see discussion around (H) in the main text) determined by $n=6$ coefficients; given the symmetry of the system, they reduce to $3$ since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
$\{h_{i,\zeta}\}_{\zeta=x,y,z}=\{h_{\zeta}\}_{\zeta=x,y,z}=\{p_{1},p_{2},-p_{1}-p_{2}\}$ (18)
A expected, we find $n-D=2$ parameters $p_{1},p_{2}$ which are used to optimise the shadow norm in (12). For the particular solution (see discussion around (P)) of $\sum_{i=\pm}\sum_{\zeta}q_{i,\zeta}E^{i}_{\zeta}=A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ we can obtain the particular coefficients
$q_{\pm,\zeta}=a\pm 3\zeta$ (19)
where $\zeta=x,y,z$. In order to find the expression for the generic projector |$\outerproduct{\psi}{\psi}(\theta,\phi)$, one can simply rely on the parametrisation $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$. This particular solution corresponds to the coefficients found with the estimators obtained with the standard inversion map, as introduced in Appendix B.0.1.
Appendix B Shadow norm expression for local Clifford measurements
In this section, we detail the explicit expressions for the shadow norm obtained using the standard estimators (13), as introduced in [2].
B.0.1 Shadow norm for local Pauli measurements
The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$. For a generic single-qubit observable $A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ the shadow norm in (10) simply evaluates to
$\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (20)
Considering the particular case of projectors on the Bloch sphere, with the parameterization $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$, eq. (20) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$.
B.0.2 Shadow norm for local Pauli measurements on the plane
In this simplified case, the set of POVMs spend a subspace on the equatorial plane of the Bloch sphere, i.e., $\mathcal{V}_{D=3}=\mathrm{span}\{\mathbbm{1},X,Y\}$. The generic target observable can be thus represented as
$A=\spmqty{undef}.$ ) . (21)
As a reference, we use the ansatz from [2] to find the inversion map in the form
$\hat{\eta}_{k}=\mathcal{M}^{-1}[E_{k}]=\mathcal{N}(\lambda E_{k}-\mu\,\mbox{Tr}\left[{E_{k}}\right]\mathbbm{1}_{2}).$ (22)
A simple inspection shows $\mathcal{N}=2$, $\lambda=2$ and $\mu=\frac{1}{2}$. The shadow norm of the generic observable results as
$\|A\|^{2}_{\text{sh}}=a^{2}+2(x^{2}+y^{2})+2|a|\sqrt{x^{2}+y^{2}}.$ (23)
For the case of projectors we have $x\to\cos{\phi}/2$ and $y\to\sin{\phi}/2$, thus we get |$\|\outerproduct{\psi(\phi)}{\psi(\phi)}\|^{2}_{\text{sh}}\equiv\frac{5}{4}$.

### v2

@@
 Introduction
-Quantum state tomography with classical shadows [1, 2], initially motivated by the seminal work of Aaronson [3], represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems.
-This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications [4, 5, 6, 7], which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead[8, 9, 10, 11, 12, 13, 14, 15, 16].
-At the core of it lies the concept of constructing these aforementioned shadows; efficiently representable compressed classical descriptions of quantum states.
+Quantum state tomography with classical shadows Huang et al.
+(2020), initially motivated by the seminal work of Aaronson Aaronson (2017), represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems.
+This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications Hu et al.
+(2023); Elben et al.
+(2020); Huang et al.
+(2021); Neven et al.
+(2021); Crawford et al.
+(2021); Chen et al.
+(2021); Hadfield et al.
+(2022); Zhao et al.
+(2021); Seif et al.
+(2023); Wan et al.
+(2023); Levy et al.
+(2024); Bu et al.
+(2024); Coopmans et al.
+(2023); Bertoni et al.
+(2023); Koh and Grewal (2022); McGinley and Fava (2023); Kunjummen et al.
+(2023); Iosue et al.
+(2024); Nakaji et al.
+(2023); Feldman et al.
+(2022); Wyderka et al.
+(2023), which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead Sack et al.
+(2022); Anshu et al.
+(2021); Huang et al.
+(2023); Nguyen et al.
+(2022); Zhang et al.
+(2021); Struchalin et al.
+(2021); Stilck França et al.
+(2024); Ippoliti et al.
+(2023); Morris et al.
+(2022); Glos et al.
+(2022); Stricker et al.
+(2022); Vermersch et al.
+(2023).
+Naturally, the main focus has been the ’quantum’ element of the procedure as this is where most practical efforts are stymied, either in gate or sample complexity.
+Other than machine learning Huang et al.
+(2023, 2022) and adaptive techniques García-Pérez et al.
+(2021); Glos et al.
+(2022), little thought has been put towards explicit optimisation of the classical postprocessing with it often being treated as a mostly fixed step.
+At the core of shadow tomography lies the concept of constructing these aforementioned shadows; efficiently representable compressed classical descriptions of quantum states.
 These shadows are generated through a process involving random measurements on copies of a state, followed by classical processing to reconstruct a succinct representation of the state, the ’shadow‘.
 These classical shadows, though not providing a complete description of the quantum state, contain enough information to estimate a wide range of properties with high accuracy when an appropriate POVM is used.
-An archetypal example is the set of measurements performed by random perturbations sampled from the Clifford group [17, 18] followed by a computational basis measurement
+An archetypal example is the set of measurements performed by random perturbations sampled from the Clifford group Webb (2016); Zhu (2017) followed by a computational basis measurement
 $\mathcal{C}_{N}=\{T\in SU(2^{N})|TP_{N}T^{\dagger}\in\mathcal{P}_{N}\}/U(1),$ (1)
 for all $P_{N}$, the set of Pauli operators on $N$ qubits.
-Critically for shadow tomography, we have that the Clifford group constitutes a $2,3$-design [17] i.e., it can reproduce the statistical properties of the full set of unitaries up to the second and third moment.
+Critically for shadow tomography, we have that the Clifford group constitutes a $2,3$-design Webb (2016) i.e., it can reproduce the statistical properties of the full set of unitaries up to the second and third moment.
 It is this that forms the basis upon which shadow tomography derives its impressive performance, implementing a map whose consequent inverse lies at the core of efficient estimation with classical shadows.
 If $U_{k}$ are elements of the $N$ qubit Clifford group and $\Pi_{\ell}$ are computational measurement projectors all acting on an input $N$ qubit quantum state then
-$\mathcal{M}[\rho]=\frac{1}{|\mathcal{C}_{N}|}\sum_{k=1}^{|\mathcal{C}_{N}|}\sum_{\ell=1}^{2^{N}}T_{k\ell}\rho T_{k\ell}=\lambda\rho+(1-\lambda)\frac{\mathbbm{1}}{2^{N}},$ (2)
+[…]
 for $T_{k\ell}=U_{k}^{\dagger}\Pi_{\ell}U_{k}$.
-The titular inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$.
-Classical shadows great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum.
+The inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$.
+Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum.
 This rapid convergence to a known configuration means the linear map of (2) may be readily inverted to recover the original state $\rho$.
-As shown in [2], the inverse map of (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
+As shown in Huang et al.
+(2020), the inverse map of (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
 Figure 1: (a) Classical shadow tomography with optimised shadow inversion.
@@
 Nevertheless, what is often neglected in studies is the fact that such a map is not unique, i.e., as long as the POVM used is overcomplete (in the operator, Hilbert-Schmidt space), an additional set of free parameters appear in the inversion process which may be optimised to improve the post-processing performance of the estimation procedure.
-This is precisely the main task of this work – we will show the general inverse map contains a free homogeneous component $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ that may be further optimised.
+Surprisingly, this has been only recently recognised in the literature Innocenti et al.
+(2023) by employing the techniques of measurement dual frames.
+In this work, we demonstrate that such additional degrees of freedom reside in the homogeneous space, i.e.
+the general inverse map contains a free homogeneous component $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ which can be further optimized.
 Our goal is to pursue such an optimisation by considering a hereto before missed resource in the tomography procedure that appears most vividly in classical shadows and the post-processing thereof.
@@
 Shadow map inversion
-We begin by describing a generalised procedure, depicted in Figure 1(a) fulfilling the random measurement scheme with a fixed POVM $\{E_{k}\}_{k=1}^{n}$ [19, 20] on a $d$-dimensional system.
+We begin by describing a generalised procedure Morris and Dakić (2020); Nguyen et al.
+(2022), depicted in Figure 1(a) fulfilling the random measurement scheme with a fixed POVM $\{E_{k}\}_{k=1}^{n}$ on a $d$-dimensional system.
 The POVM elements sample randomly an unknown state $\rho$ with probability $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$.
@@
 each outcome $E_{k}$ the corresponding matrix $\hat{\eta}_{k}$, which represents a single-shot approximation of $\bar{\rho}$.
-They can be calculated via an inversion map $\hat{\eta_{k}}=\mathcal{M}^{-1}(E_{k})$ [2, 20] which we will show is not unique, i.e.
-$\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ with the homogeneous component containing auxiliary parameters that may be used as a free resource in post-processing.
+They can be calculated via an inversion map $\hat{\eta_{k}}=\mathcal{M}^{-1}(E_{k})$ Huang et al.
+(2020); Nguyen et al.
+(2022) which we will show is not unique, i.e.
+$\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ with the latter (homogeneous) component containing auxiliary parameters that may be used as a free resource in post-processing.
 The overall classical shadow itself is obtained by averaging over the estimators obtained from $M$ random measurements: $M^{-1}\sum_{s}^{M}\hat{\eta}^{(s)}$.
@@
 By substituting $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{AE_{k}}\right]$ into the previous equation we find
-$\sum_{k}a_{k}^{(p)}E_{k}=A,$ (P)
-$\sum_{k}a_{k}^{(h)}E_{k}=0.$ (H)
+$\sum_{k}a_{k}^{(p)}E_{k}=A,$ (10)
+$\sum_{k}a_{k}^{(h)}E_{k}=0.$ (11)
 The set of solutions to the latter (homogeneous) equation defines an auxiliary optimisation space of dimension $n-D$ independent of the target observable (fixed by the choice of POVM).
@@
 The contour plot shows the minimal shadow norm (13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$).
-Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map (13) found in [2] (b) Optimal free parameter for plane projectors.
+Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map (15) found in Huang et al.
+(2020) (b) Optimal free parameter for plane projectors.
 The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$).
@@
 Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy.
-It is customary in classical shadow tomography to refer to variance as figure of merit [2].
+It is customary in classical shadow tomography to refer to variance as figure of merit Huang et al.
+(2020).
 In particular, we consider the state-independent upper bound on variance, defined as the shadow norm, which only depends on the decomposition of the particular observable:
@@
 This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably.
-This is akin to the original upper bound on shadow norm and thence sampling complexity found in [2], where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{shadow}})$ ).
+This is akin to the original upper bound on shadow norm and thence sampling complexity found in Huang et al.
+(2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{shadow}})$ ).
 Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator.
@@
 Let us now discuss the estimation of product observables $A^{(N)}=A_{1}\otimes A_{2}\ldots\otimes A_{N}$ via local POVMs, i.e., of the type $E_{k_{1}}\otimes\dots\otimes E_{k_{n}}$.
-This case is particularly interesting for near-term applications, such as the VQE algorithm [21, 22].
+This case is particularly interesting for near-term applications, such as the VQE algorithm Peruzzo et al.
+(2014); Tilly et al.
+(2022).
 In such a scenario, it is easy to show that the shadow norm in Eq.
-(10) has a product form (see also [20]) and consequently Eq.
+(12) has a product form (see also Nguyen et al.
+(2022)) and consequently Eq.
 (13) becomes
@@
 $\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{3}\right).$ (15)
-found in [2].
+found in Huang et al.
+(2020).
 Now we show how to improve this bound by our techniques.
 The POVMs $E^{\pm}_{\zeta}$ span the entire HS space.
-We have $n=6$ and $D=4$, thus the equation (H) gives two free optimisation parameters for each qubit (details can be found in appendix A).
+We have $n=6$ and $D=4$, thus the equation (11) gives two free optimisation parameters for each qubit See supplementary material at [URL from publisher] for information concerning estimator explicit parameter dependence .
 In Fig.2(a), we present the minimal upper bound as a function of angles $(\theta,\phi)$, which is contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$.
@@
 The target observables are tensor products of a projector on a single qubit, parameterised by a single angle: $A=\outerproduct{\psi(\phi)}{\psi(\phi)}$ | with $|{\psi(\phi)}\rangle=(|{0}\rangle+e^{i\phi}|{1}\rangle)\sqrt{2}$.
-As before, we follow the ansatz given in [2] to get the particular solution for the inversion map and estimators
+As before, we follow the ansatz given in Huang et al.
+(2020) to get the particular solution for the inversion map and estimators
 $\hat{\eta}^{\pm}_{\zeta}=4E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}.$ (16)
 From here we obtain the reference bound for the shadow norm of $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\equiv(5/4)^{N}$.
-We reference to Appendix B.0.2 for more precise remarks on how we obtain this result.
+We reference to Supplementary Information See supplementary material at [URL from publisher] for information concerning shadow norm expression for local Clifford measurements for more precise remarks on how we obtain this result.
 Now we turn to our optimisation technique.
@@
 Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimisation will be the topic of future work.
+Note - In preparation of this manuscript, we became aware of two other works Malmi et al.
+(2024); Fischer et al.
+(2024) considering a similar problem.
+The authors consider a similar optimisation of the post-processing stage in estimation tasks, albeit referring to different frameworks.
 Acknowledgements
 We thank Richard Küng for his insight and excellent discussions on the subject matter.
-This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/F7115] (BeyondC).
+This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/F71] (BeyondC).
 For the purpose of open access, the author(s) has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
 Appendix A Parameter dependence for estimators
-The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$, admits a homogeneous combination $\sum_{i=\pm}\sum_{\zeta}h_{i,\zeta}E^{i}_{\zeta}=0$ (see discussion around (H) in the main text) determined by $n=6$ coefficients; given the symmetry of the system, they reduce to $3$ since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
+The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$, admits a homogeneous combination $\sum_{i=\pm}\sum_{\zeta}h_{i,\zeta}E^{i}_{\zeta}=0$ (see discussion around (11) in the main text) determined by $n=6$ coefficients; given the symmetry of the system, they reduce to $3$ since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
 $\{h_{i,\zeta}\}_{\zeta=x,y,z}=\{h_{\zeta}\}_{\zeta=x,y,z}=\{p_{1},p_{2},-p_{1}-p_{2}\}$ (20)
 A expected, we find $n-D=2$ parameters $p_{1},p_{2}$ which are used to optimise the shadow norm in (14).
-For the particular solution (see discussion around (P)) of $\sum_{i=\pm}\sum_{\zeta}q_{i,\zeta}E^{i}_{\zeta}=A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ we can obtain the particular coefficients
+For the particular solution (see discussion around (10)) of $\sum_{i=\pm}\sum_{\zeta}q_{i,\zeta}E^{i}_{\zeta}=A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ we can obtain the particular coefficients
 $q_{\pm,\zeta}=a\pm 3\zeta$ (21)
@@
 In order to find the expression for the generic projector |$\outerproduct{\psi}{\psi}(\theta,\phi)$, one can simply rely on the parametrisation $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$.
-This particular solution corresponds to the coefficients found with the estimators obtained with the standard inversion map, as introduced in Appendix B.0.1.
+This particular solution corresponds to the coefficients found with the estimators obtained with the standard inversion map, as introduced in B.0.1.
 Appendix B Shadow norm expression for local Clifford measurements
-In this section, we detail the explicit expressions for the shadow norm obtained using the standard estimators (13), as introduced in [2].
+In this section, we detail the explicit expressions for the shadow norm obtained using the standard estimators (15), as introduced in the seminal work on Classical Shadows.
 B.0.1 Shadow norm for local Pauli measurements
@@
 $\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (22)
-Considering the particular case of projectors on the Bloch sphere, with the parameterization $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$, eq.
-(20) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$.
+Considering the particular case of projectors on the Bloch sphere, with the parameterization $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$, (22) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$.
 B.0.2 Shadow norm for local Pauli measurements on the plane
@@
 $A=\spmqty{undef}.$ ) . (23)
-As a reference, we use the ansatz from [2] to find the inversion map in the form
+As a reference, we use the ansatz from Huang et al. (2020) to find the inversion map in the form
 $\hat{\eta}_{k}=\mathcal{M}^{-1}[E_{k}]=\mathcal{N}(\lambda E_{k}-\mu\,\mbox{Tr}\left[{E_{k}}\right]\mathbbm{1}_{2}).$ (24)

### v3

@@
 (2020); Huang et al.
-(2021); Neven et al.
+(2021a); Neven et al.
 (2021); Crawford et al.
@@
 Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum.
-This rapid convergence to a known configuration means the linear map of (2) may be readily inverted to recover the original state $\rho$.
+This rapid convergence to a known configuration means the linear map of Eq.
+(2) may be readily inverted to recover the original state $\rho$.
 As shown in Huang et al.
-(2020), the inverse map of (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
+(2020), the inverse map of Eq.
+(2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
 Figure 1: (a) Classical shadow tomography with optimised shadow inversion.
@@
 Assuming that all samples are i.i.d., the average value $\expectationvalue{\hat{\eta}}$ is immediately
-$\expectationvalue{\hat{\eta}}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}\overset{!}{=}\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
+$\expectationvalue{\hat{\eta}}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}=\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
 This is the requirement on the estimator $\hat{\eta}$ such that the statistics of the shadow mimics, in expectation, that of the actual unknown state.
-Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, any solution to Eq.
+With Eq.
+(3), the task of finding an inversion map can be made equivalent to identifying a dual frame to the POVM in space of operators Innocenti et al.
+(2023).
+Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, the solution to Eq.
 (3) is not unique.
@@
 To achieve this, we now show a constructive method to recover families of equivalent classical shadows in terms of the POVM elements and their free parameters.
-We rely on an equivalent problem to (3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
-$\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]\overset{!}{=}\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
+We rely on an equivalent problem to Eq.
+(3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
+$\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]=\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
 It is instructive to express the operators as vectors in $\mathcal{V}_{D}$, i.e., for an operator $X\in\mathcal{V}_{D}$, we shall use the ket notation $|{X}\rangle\!\rangle$.
 We also identify an orthonormal basis $\beta_{D}=\{|{B_{j}}\rangle\!\rangle\}_{j=1}^{D}$ in $\mathcal{V}_{D}$.
-The expectation value from (4) can be expressed as
-$\expectationvalue{\,a\,}=\sum_{k=1}^{n}\langle\!\langle{A^{\dagger}}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|{\rho}\rangle\!\rangle\overset{!}{=}\langle\!\langle{A^{\dagger}}|{\bar{\rho}}\rangle\!\rangle$ (5)
-where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$.
+The expectation value from Eq. (4) can be expressed as
+$\expectationvalue{\,a\,}=\sum_{k=1}^{n}\langle\!\langle{A^{\dagger}}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|{\rho}\rangle\!\rangle=\langle\!\langle{A^{\dagger}}|{\bar{\rho}}\rangle\!\rangle$ (5)
+where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$ is the inner product in $\mathcal{V}_{D}$.
 To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$
@@
 which can be solved for $|{\eta_{k}}\rangle\!\rangle$ using $E_{k}^{\dagger}=E_{k}$. The last equation expressed in the basis $\beta_{D}$ reads
-$\sum_{k=1}^{n}\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|{B_{t}}\rangle\!\rangle=\sum_{k=1}^{n}L_{sk}R_{tk}^{*}\overset{!}{=}\delta_{st},$ (7)
+$\sum_{k=1}^{n}\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|{B_{t}}\rangle\!\rangle=\sum_{k=1}^{n}L_{sk}R_{tk}^{*}=\delta_{st},$ (7)
 with $L_{sk}=\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle$ and $R_{tk}=\langle\!\langle{B_{t}}|{E_{k}}\rangle\!\rangle$.
@@
 We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero-matrix.
-Having this, a simple exercise shows $L=W(\Sigma^{-1}|H)U^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix.
-Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=W(\Sigma^{-1}|\hat{0})U^{\dagger}$ (particular) and $L^{(h)}=W(\hat{0}|H)U^{\dagger}$ (homogeneous).
+Having this, a simple exercise shows $L=U(\Sigma^{-1}|H)W^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix.
+Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=U(\Sigma^{-1}|\hat{0})W^{\dagger}$ (particular) and $L^{(h)}=U(\hat{0}|H)W^{\dagger}$ (homogeneous).
 This results in a family of equivalent estimators
@@
 As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimisation in post-processing.
-We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$ where the homogeneous coefficients do not influence expectation values or any other linear combination of classical shadows
+We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$.
+The homogeneous coefficients, from the definition of $\hat{\eta}_{k}^{(h)}$ in Eq.
+(8), do not influence expectation values or any other linear combination of classical shadows
 $\expectationvalue{\,a\,}=\sum_{k=1}^{n}p_{k}a_{k}=\sum_{k=1}^{n}p_{k}\left(a_{k}^{(p)}+a_{k}^{(h)}\right)=\sum_{k=1}^{n}p_{k}a_{k}^{(p)}.$ (9)
-By substituting $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{AE_{k}}\right]$ into the previous equation we find
+By substituting $p_{k}=\mbox{Tr}\left[{\rho\,E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\,\rho}\right]$ into the previous equation we find
 $\sum_{k}\,a_{k}^{(p)}\,E_{k}=A,$ (10)
@@
 Figure 2: (a) Optimal free parameter for projectors on the Bloch sphere.
-The contour plot shows the minimal shadow norm (13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$).
-Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map (15) found in Huang et al.
+The contour plot shows the minimal shadow norm Eq.
+(13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$).
+Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map Eq.
+(15) found in Huang et al.
 (2020) (b) Optimal free parameter for plane projectors.
 The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$).
-This enhanced norm is always lower or equal to the standard shadow norm |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\leq 5/4$ (see main text).
-In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis.
-The bound corresponds to the combinations which present optimal variance |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
+This enhanced norm takes values in the range $41\leq\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\leq 5/$ 4 (see main text).
+In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis, the bound achieves optimal value |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
 Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy.
@@
 (2020).
-In particular, we consider the state-independent upper bound on variance, defined as the shadow norm, which only depends on the decomposition of the particular observable:
-$\mbox{Var}\left[{a}\right]\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{shadow}}.$ (12)
+In particular, we consider the state-independent upper bound on variance, called shadow norm $\|A\|^{2}_{\text{sh}}$, defined as:
+$\begin{split}\mbox{Var}\left[{a}\right]&=\mathbb{E}[a^{2}]-(\mathbb{E}[a])^{2}\leq\mathbb{E}[a^{2}]=\sum_{k=1}^{n}p_{k}|a_{k}|^{2}\\ &\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{sh}}.\end{split}$ (12)
+Considering the upper bound to study sample complexity is reasonable since, for bounded observables, the quantity $(\mathbb{E}[a])^{2}$ is also bounded, therefore the relevant contribution is brought by the first term $\mathbb{E}[a^{2}]$.
 This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably.
 This is akin to the original upper bound on shadow norm and thence sampling complexity found in Huang et al.
-(2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{shadow}})$ ).
+(2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{sh}})$ ).
 Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator.
@@
 The existence of a minimum is guaranteed through the positive semidefiniteness of the variance operator.
-We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in (13) yields exceptional gains for relatively minuscule cost.
+We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in Eq.
+(13) yields exceptional gains for relatively minuscule cost.
 Product observables and shadow optimisation
@@
 $\|A^{(N)}\|^{2}_{\text{opt}}=\prod_{j=1}^{N}\|A_{j}\|^{2}_{\text{opt}}$ (14)
-The optimisation in Eq.
-(12) appears efficient, given the total number of free parameters scales linearly with the number of subsystems $N$.
+Since we are considering local measurements on independent subsystems, each term $\|A_{j}\|^{2}_{\text{opt}}$ can be optimised locally, implying the optimisation in Eq.
+(14) is efficient overall.
 We see that even minimal deviation from the minimum in local shadow norms $\|A_{j}\|^{2}_{\text{opt}}$ may accumulate exponentially fast with $N$.
@@
 The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterised on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$.
-The measurement scheme is described by the POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$.
-As reference, we consider the bound $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{shadow}}\equiv(3/2)^{N}$ obtained using the estimators
-$\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{3}\right).$ (15)
-found in Huang et al.
-(2020).
-Now we show how to improve this bound by our techniques.
-The POVMs $E^{\pm}_{\zeta}$ span the entire HS space.
-We have $n=6$ and $D=4$, thus the equation (11) gives two free optimisation parameters for each qubit See supplementary material at [URL from publisher] for information concerning estimator explicit parameter dependence .
-In Fig.2(a), we present the minimal upper bound as a function of angles $(\theta,\phi)$, which is contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$.
-Since the overall shadow norm (14) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
+The measurement scheme is described by the POVM of normalised projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{6}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$.
+Using the estimators
+$\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}\right).$ (15)
+found in Huang et al. (2020), the shadow norm in Eq. (12) simply evaluates for the generic single-qubit observable $A=a\cdot\mathbbm{1}+x\cdot\sigma_{x}+y\cdot\sigma_{y}+z\cdot\sigma_{z}$ to
+$\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (16)
+Considering the particular case of projectors on the Bloch sphere, with the parameterisation
+$\begin{split}a\to\frac{1}{2},\quad\quad x\to\frac{1}{2}\sin\theta\cos\phi,\\ y\to\frac{1}{2}\sin\theta\sin\phi,\quad\quad z\to\frac{1}{2}\cos\theta,\end{split}$ (17)
+and Eq.
+(16) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$.
+Now we show how to improve this reference bound by our techniques.
+The POVMs $E^{\pm}_{\zeta}$ span the entire HS space and we may determine easily a homogeneous combination of the effects as in Eq.
+(11) given $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$.
+Exploiting the symmetry of the system, they reduce to three parameters since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
+$\{h_{\zeta}\}_{\zeta=x,y,z}=\{p_{1},p_{2},-p_{1}-p_{2}\}$ (18)
+As expected, we find $n-D=2$ parameters $p_{1},p_{2}$.
+Optimising over these parameters, we present the minimal upper bound as a function of angles $(\theta,\phi)$ in Fig.2(a).
+The optimised results are contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$.
+Since the overall shadow norm Eq.
+(14) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
+Estimation with constrained POVMs
+In the previous section we have studied examples of overcomplete POVMs in the entire HS space, i.e.
+with $D=d^{2}$.
+Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an over-complete basis (within $\mathcal{V}_{D}$).
+Note that the standard inversion procedure Huang et al.
+(2020); Nguyen et al.
+(2022) is defined on the full HS space and does not directly apply in this instance.
+Consider for instance the POVM constituted by the projectors of the X and Y Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the ”equator” on the Bloch sphere.
+With a straightforward application of our procedure Eq.
+(8), we obtain the following particular solution
+$\hat{\eta}^{\pm}_{\zeta}=4E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}.$ (19)
+The homogeneous term is added by solving $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$, where $\{h_{\zeta}\}_{\zeta=x,y}=\{p,-p\}$ and $p$ is a free parameter .
+For the generic planar observable (within span of $\{E_{\zeta}^{\pm}\}$)
+$A=\matrixquantity(a&x+iy\\ x-iy&a),$ (20)
+we find the shadow norm
+$\|A\|^{2}_{\text{sh}}(p)$ $=(a-x)^{2}+2(x^{2}+y^{2})+B(p)+\sqrt{C(p)},$ (21)
+$B(p)$ $=\frac{1}{8}(2(p_{r}^{2}+p_{i}^{2})a^{2}+(p_{r}(8+p_{r})+p_{i}^{2})x^{2}$
+$+(p_{r}(p_{r}-8)+p_{i}^{2})y^{2}),$
+$C(p)$ $=x^{2}(a(8+p_{r}(6+p_{r})+p_{i}^{2})-2(4+p_{r})x)^{2}$
+$+(a(8+(p_{r}-6)p_{r}+p_{i}^{2})+2(p_{r}-4)x)^{2}y^{2},$
+[…]
+We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies Huang et al. (2020).
+Optimal estimation of planar Paulis
+When considering as target observable the planar Pauli,
+$a\to 0,\quad x\to\cos\varphi,\quad y\to\sin\varphi,$ (22)
+the optimised shadow norm from Eq.
+(21) results in $\|A\|^{2}_{\text{opt}}=2$.
+An immediate result is that the expected number of resources for product of planar Paulis scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements Huang et al.
+(2020), $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$.
+Surprisingly, we are able to match the optimal scaling Huang et al.
+(2021b) while still relying only on local single-qubit measurements.
+Complete versus overcomplete POVMs
+In the context of the previous example, it is interesting to compare the performance of complete versus overcomplete POVM.
+For this purpose, consider the example of a triangle POVM on the plane, defined as the normalised projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
+$|{\psi_{k}}\rangle=\dfrac{|{0}\rangle+e^{i\frac{k}{3}2\pi}|{1}\rangle}{\sqrt{2}}\quad k=0,1,2$ (23)
+Simple inversion of this complete basis brings us to estimators
+$\hat{\eta}_{k}=3E_{k}-\frac{3}{4}\mbox{Tr}\left[{E_{k}}\right]\mathbbm{1}_{2}.$ (24)
+Since the POVM is complete, there is no free parameter dependence.
+From here, we can compute the shadow norm as defined in Eq.(12).
+For a planar Pauli as target observable, as in Eq.
+(22), the shadow norm is $\|A\|^{2}_{\text{sh}}=3,\,\forall\varphi$, which leads to an expected overhead of $3^{N}$ on the scaling of resources.
+From the previous example, the same observables can be optimally estimated with an overhead of $2^{N}$.
+We have therefore found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
 Locally efficient fidelity estimation implies globally efficient
-We now present a configuration in which the optimisation of the inversion map not only improves the standard bound, but we are actually able to achieve efficiency, i.e.
-$\|A\|^{2}_{\text{opt}}\equiv 1$.
-We consider the POVM constituted by the projectors of the X and Y Paulis, which spans the ”equator” on the Bloch sphere.
-The target observables are tensor products of a projector on a single qubit, parameterised by a single angle: $A=\outerproduct{\psi(\phi)}{\psi(\phi)}$ | with $|{\psi(\phi)}\rangle=(|{0}\rangle+e^{i\phi}|{1}\rangle)\sqrt{2}$.
-As before, we follow the ansatz given in Huang et al.
-(2020) to get the particular solution for the inversion map and estimators
-$\hat{\eta}^{\pm}_{\zeta}=4E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}.$ (16)
-From here we obtain the reference bound for the shadow norm of $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\equiv(5/4)^{N}$.
-We reference to Supplementary Information See supplementary material at [URL from publisher] for information concerning shadow norm expression for local Clifford measurements for more precise remarks on how we obtain this result.
-Now we turn to our optimisation technique.
-We have the POVM with $n=4$ elements spanning a subspace of dimension $D=3$, therefore the generic estimators obtained from our procedure depend on a single free parameter $p$.
-The shadow norm of single-qubit projector depends on this free parameter and on the angle parameter $\phi$:
+We now present a configuration in which the optimisation of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e.
+$\|A\|^{2}_{\text{opt}}=1$.
+To show this, we choose as target observables tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ | where each single-qubit state is parameterised as: $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
+Using the change of coordinates
+$a\to\frac{1}{2},\quad x\to\frac{1}{2}\cos\phi,\quad y\to\frac{1}{2}\sin\phi,$ (25)
+we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$:
 |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}=\frac{1}{2}$ $(1+p_{i}^{2}+B(\phi,p_{r})+$ (26)
@@
 $\quad+2p_{r}(p_{r}-1)$
-$C(\phi,p_{r})=(4$ $)\displaystyle p_{r}-1)\cos{\phi}+2p_{r}\cos{2\phi}$ ) (19)
+$C(\phi,p_{r})=(4$ $)\displaystyle p_{r}-1)\cos{\phi}+2p_{r}\cos{2\phi}$ )
 $\quad+\cos{3\phi}+2+2p_{r}(p_{r}-1),$
 […]
-In general, the dependency on $\phi$ remains explicit, changing the expected bound based on the projector to be estimated.
-In Fig.2(b) we present the optimal value of the shadow norm for each parameter $\phi\in[0,\pi]$.
-We note that for each value of $\phi$ the optimised norm is |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{opt}}\leq 5/4$.
-Moreover, it can be seen that for specific angles, the bound is exactly $\|E_{\zeta}^{\pm}\|^{2}_{\text{opt}}\equiv 1$., i.e., for the eigenprojectors of $X$ and $Y$ Paulis, corresponding to the POVM elements themselves.
-This implies efficient estimation: given the product form of the overall shadow norm (14), this scaling represents not only an exponential improvement compared to $(5/4)^{N}$, as for other projectors, but also an expected constant scaling of the sample size with the number of qubits.
-This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $2^{N}$ projections simultaneously for complementary observables.
-Previously this result has been found only via global measurement strategies relying on Clifford unitaries consisting of many joint operations, acting on all $N$ qubits at the same time; we have shown instead that by introducing the initial assumption over the space of observables, this could be achieved using only local single-qubit measurements.
+This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $4^{N}$ projections simultaneously for complementary observables.
+Previously this result has been found in Huang et al.
+(2020) only via global measurement strategies relying on $N-$qubit Clifford unitaries.
+We have shown instead that this is achieved using exclusively local measurements.
 Discussion
@@
 For the purpose of open access, the author(s) has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
-Appendix A Parameter dependence for estimators
-The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$, admits a homogeneous combination $\sum_{i=\pm}\sum_{\zeta}h_{i,\zeta}E^{i}_{\zeta}=0$ (see discussion around (11) in the main text) determined by $n=6$ coefficients; given the symmetry of the system, they reduce to $3$ since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
-$\{h_{i,\zeta}\}_{\zeta=x,y,z}=\{h_{\zeta}\}_{\zeta=x,y,z}=\{p_{1},p_{2},-p_{1}-p_{2}\}$ (20)
-A expected, we find $n-D=2$ parameters $p_{1},p_{2}$ which are used to optimise the shadow norm in (14).
-For the particular solution (see discussion around (10)) of $\sum_{i=\pm}\sum_{\zeta}q_{i,\zeta}E^{i}_{\zeta}=A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ we can obtain the particular coefficients
-$q_{\pm,\zeta}=a\pm 3\zeta$ (21)
-where $\zeta=x,y,z$.
-In order to find the expression for the generic projector |$\outerproduct{\psi}{\psi}(\theta,\phi)$, one can simply rely on the parametrisation $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$.
-This particular solution corresponds to the coefficients found with the estimators obtained with the standard inversion map, as introduced in B.0.1.
-Appendix B Shadow norm expression for local Clifford measurements
-In this section, we detail the explicit expressions for the shadow norm obtained using the standard estimators (15), as introduced in the seminal work on Classical Shadows.
-B.0.1 Shadow norm for local Pauli measurements
-The POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}\left(\mathbbm{1}\pm\sigma_{\zeta}\right)$, where $\zeta=x,y,z$.
-For a generic single-qubit observable $A=a\mathbbm{1}+\sum_{\zeta=x,y,z}\zeta\sigma_{\zeta}$ the shadow norm in (12) simply evaluates to
-$\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (22)
-Considering the particular case of projectors on the Bloch sphere, with the parameterization $a\to\frac{1}{2}$, $x\to\frac{1}{2}\sin\theta\cos\phi$, $y\to\frac{1}{2}\sin\theta\sin\phi$, $z\to\frac{1}{2}\cos\theta$, (22) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$.
-B.0.2 Shadow norm for local Pauli measurements on the plane
-In this simplified case, the set of POVMs spend a subspace on the equatorial plane of the Bloch sphere, i.e., $\mathcal{V}_{D=3}=\mathrm{span}\{\mathbbm{1},X,Y\}$.
-The generic target observable can be thus represented as
-$A=\spmqty{undef}.$ ) . (23)
-As a reference, we use the ansatz from Huang et al. (2020) to find the inversion map in the form
-$\hat{\eta}_{k}=\mathcal{M}^{-1}[E_{k}]=\mathcal{N}(\lambda E_{k}-\mu\,\mbox{Tr}\left[{E_{k}}\right]\mathbbm{1}_{2}).$ (24)
-A simple inspection shows $\mathcal{N}=2$, $\lambda=2$ and $\mu=\frac{1}{2}$. The shadow norm of the generic observable results as
-$\|A\|^{2}_{\text{sh}}=a^{2}+2(x^{2}+y^{2})+2|a|\sqrt{x^{2}+y^{2}}.$ (25)
-For the case of projectors we have $x\to\cos{\phi}/2$ and $y\to\sin{\phi}/2$, thus we get |$\|\outerproduct{\psi(\phi)}{\psi(\phi)}\|^{2}_{\text{sh}}\equiv\frac{5}{4}$.

### v4

@@
-Optimising quantum tomography via shadow inversion
+Optimizing quantum tomography via shadow inversion
 Abstract
-In quantum information theory, the accurate estimation of observables is pivotal for quantum information processing, playing a crucial role in compute and communication protocols.
-This work introduces a novel technique for estimating such objects, leveraging an underutilised resource in the inversion map of classical shadows that greatly refines the estimation cost of target observables without incurring any additional overhead.
-A generalised framework for computing and optimising additional degrees of freedom in the homogeneous space of the shadow inversion is given that may be adapted to a variety of near-term problems.
-In the special case of local measurement strategies we show feasible optimisation leading to an exponential separation in sample complexity versus the standard approach and in an exceptional case we give non-trivial examples of optimised post-processing for local measurements, achieving the same efficiency as the global Cliffords shadows.
+In quantum information theory, the accurate estimation of observables is pivotal for quantum information processing, playing a crucial role in computational and communication protocols.
+This work introduces a technique for estimating such objects, leveraging an underutilized resource in the inversion map of classical shadows that greatly refines the estimation cost of target observables without incurring any additional overhead.
+A generalized framework for computing and optimizing additional degrees of freedom in the homogeneous space of the shadow inversion is given that may be adapted to a variety of near-term problems.
+In the special case of local measurement strategies we show feasible optimization leading to an exponential separation in sample complexity versus the standard approach and, in an exceptional case, we give nontrivial examples of optimized postprocessing for local measurements, achieving the same efficiency as the global Cliffords shadows.
 Introduction
-Quantum state tomography with classical shadows Huang et al.
-(2020), initially motivated by the seminal work of Aaronson Aaronson (2017), represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems.
-This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications Hu et al.
-(2023); Elben et al.
-(2020); Huang et al.
-(2021a); Neven et al.
-(2021); Crawford et al.
-(2021); Chen et al.
-(2021); Hadfield et al.
-(2022); Zhao et al.
-(2021); Seif et al.
-(2023); Wan et al.
-(2023); Levy et al.
-(2024); Bu et al.
-(2024); Coopmans et al.
-(2023); Bertoni et al.
-(2023); Koh and Grewal (2022); McGinley and Fava (2023); Kunjummen et al.
-(2023); Iosue et al.
-(2024); Nakaji et al.
-(2023); Feldman et al.
-(2022); Wyderka et al.
-(2023), which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead Sack et al.
-(2022); Anshu et al.
-(2021); Huang et al.
-(2023); Nguyen et al.
-(2022); Zhang et al.
-(2021); Struchalin et al.
-(2021); Stilck França et al.
-(2024); Ippoliti et al.
-(2023); Morris et al.
-(2022); Glos et al.
-(2022); Stricker et al.
-(2022); Vermersch et al.
-(2023).
-Naturally, the main focus has been the ’quantum’ element of the procedure as this is where most practical efforts are stymied, either in gate or sample complexity.
-Other than machine learning Huang et al.
-(2023, 2022) and adaptive techniques García-Pérez et al.
-(2021); Glos et al.
-(2022), little thought has been put towards explicit optimisation of the classical postprocessing with it often being treated as a mostly fixed step.
-At the core of shadow tomography lies the concept of constructing these aforementioned shadows; efficiently representable compressed classical descriptions of quantum states.
-These shadows are generated through a process involving random measurements on copies of a state, followed by classical processing to reconstruct a succinct representation of the state, the ’shadow‘.
-These classical shadows, though not providing a complete description of the quantum state, contain enough information to estimate a wide range of properties with high accuracy when an appropriate POVM is used.
-An archetypal example is the set of measurements performed by random perturbations sampled from the Clifford group Webb (2016); Zhu (2017) followed by a computational basis measurement
+Quantum state tomography with classical shadows [1], initially motivated by the seminal work of Aaronson [2], represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterization of quantum systems.
+This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography, has since seen a plethora of modifications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
+Naturally, the main focus has been the “quantum” element of the procedure as this is where most practical efforts are stymied, either in gate or sample complexity.
+Other than machine learning [26, 36] and adaptive techniques [37, 33], little thought has been put towards explicit optimization of the classical postprocessing, with it often being treated as a mostly fixed step.
+At the core of shadow tomography lies the concept of constructing these aforementioned shadows, i.e., efficiently representable compressed classical descriptions of quantum states.
+These shadows are generated through a process involving random measurements on copies of a state, followed by classical processing to reconstruct a succinct representation of the state, the “shadow”.
+These classical shadows, though not providing a complete description of the quantum state, contain enough information to estimate a wide range of properties with high accuracy when an appropriate Positive Operator-Valued Measurement (POVM) is used.
+An archetypal example is the set of measurements performed by random perturbations sampled from the Clifford group [38, 39] followed by a computational basis measurement
 $\mathcal{C}_{N}=\{T\in SU(2^{N})|TP_{N}T^{\dagger}\in\mathcal{P}_{N}\}/U(1),$ (1)
 for all $P_{N}$, the set of Pauli operators on $N$ qubits.
-Critically for shadow tomography, we have that the Clifford group constitutes a $2,3$-design Webb (2016) i.e., it can reproduce the statistical properties of the full set of unitaries up to the second and third moment.
+Critically for shadow tomography, we have that the Clifford group constitutes a $2,3$-design [38], i.e., it can reproduce the statistical properties of the full set of unitaries up to the second and third moments.
 It is this that forms the basis upon which shadow tomography derives its impressive performance, implementing a map whose consequent inverse lies at the core of efficient estimation with classical shadows.
-If $U_{k}$ are elements of the $N$ qubit Clifford group and $\Pi_{\ell}$ are computational measurement projectors all acting on an input $N$ qubit quantum state then
+If $U_{k}$ are elements of the $N-$qubit Clifford group and $\Pi_{\ell}$ are computational measurement projectors all acting on an input $N-$qubit quantum state then
 […]
@@
 The inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$.
-Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum.
+Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements is used in the above sum.
 This rapid convergence to a known configuration means the linear map of Eq.
 (2) may be readily inverted to recover the original state $\rho$.
-As shown in Huang et al.
-(2020), the inverse map of Eq.
-(2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
-Figure 1: (a) Classical shadow tomography with optimised shadow inversion.
-After a collection of $M$ POVM outcomes have been acquired from a fixed quantum state, any estimation can be improved for a dynamic choice of target observable $A$ by computing an optimal ‘inverse’ operation for that observable and POVM.
+As shown in [1], the inverse map of Eq.
+(2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarizing channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
+Figure 1: (a) Classical shadow tomography with optimized shadow inversion.
+After a collection of $M$ POVM outcomes has been acquired from a fixed quantum state, any estimation can be improved for a dynamic choice of target observable $A$ by computing an optimal “inverse” operation for that observable and POVM.
 (b) Shadow norm for different strategies.
-The shadow norm cofactor when optimised for certain target projectors, using local Pauli measurements on single qubits.
-The shadow norm for such an observable on $N$ qubits is normally $(3/2)^{N}$ but when optimised over additional degrees of freedom can be shrunk to $(1.15)^{N}$.
-Though still exponential in the number of qubits, the optimised strategy is exponentially cheaper than standard classical shadow tomography with local measurements.
+The shadow norm cofactor when optimized for certain target projectors, using local Pauli measurements on single qubits.
+The shadow norm for such an observable on $N$ qubits is normally $(3/2)^{N}$, but when optimized over additional degrees of freedom, it can be shrunk to $(1.15)^{N}$.
+Though still exponential in the number of qubits, the optimized strategy is exponentially cheaper than standard classical shadow tomography with local measurements.
 In these cases, the inversion map takes a particularly simple form, and the estimation of consumed resources (sampling complexity) critically depends on it.
-Nevertheless, what is often neglected in studies is the fact that such a map is not unique, i.e., as long as the POVM used is overcomplete (in the operator, Hilbert-Schmidt space), an additional set of free parameters appear in the inversion process which may be optimised to improve the post-processing performance of the estimation procedure.
-Surprisingly, this has been only recently recognised in the literature Innocenti et al.
-(2023) by employing the techniques of measurement dual frames.
-In this work, we demonstrate that such additional degrees of freedom reside in the homogeneous space, i.e.
-the general inverse map contains a free homogeneous component $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ which can be further optimized.
-Our goal is to pursue such an optimisation by considering a hereto before missed resource in the tomography procedure that appears most vividly in classical shadows and the post-processing thereof.
+Nevertheless, what is often neglected in studies is the fact that such a map is not unique, i.e., as long as the POVM that is used is overcomplete (in the operator, Hilbert-Schmidt space), an additional set of free parameters appears in the inversion process which may be optimized to improve the postprocessing performance of the estimation procedure.
+Surprisingly, this has been only recently recognized in the literature [40] by employing the techniques of measurement dual frames.
+In this work, we demonstrate that such additional degrees of freedom reside in the homogeneous space, i.e., the general inverse map contains a free homogeneous component, $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$, which can be further optimized.
+Our goal is to pursue such an optimization by considering a heretofore missed resource in the tomography procedure that appears most vividly in classical shadows and the postprocessing thereof.
 Though the natural choice of the inverse map (such as the one given above) is the inverse of the forward channel being applied, there is nothing that says it must be so.
 It is this inversion process, normally treated as a fixed mathematical procedure, that we aim to modify and improve upon.
-As long as the set of performed measurements is tomographically over-complete, the additional degrees of freedom may be exploited to reduce case-specific variance for a given observable.
+As long as the set of performed measurements is tomographically overcomplete, the additional degrees of freedom may be exploited to reduce case-specific variance for a given observable.
 In doing so, we will show that an exponential improvement in estimations is possible compared to the standard choice of inverse map presented in the literature.
-With this novel approach to processing acquired quantum data, we open a new route to exploit areas where advantages may be gained without incurring an unacceptable resource cost.
+With this approach to processing acquired quantum data, we open a route to exploit areas where advantages may be gained without incurring an unacceptable resource cost.
 Shadow map inversion
-We begin by describing a generalised procedure Morris and Dakić (2020); Nguyen et al.
-(2022), depicted in Figure 1(a) fulfilling the random measurement scheme with a fixed POVM $\{E_{k}\}_{k=1}^{n}$ on a $d$-dimensional system.
-The POVM elements sample randomly an unknown state $\rho$ with probability $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$.
+We begin by describing a generalized procedure [41, 27], depicted in Fig.
+1(a), fulfilling the random measurement scheme with a fixed POVM $\{E_{k}\}_{k=1}^{n}$ on a $d$-dimensional system.
+The POVM elements randomly sample an unknown state $\rho$ with probability $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$.
 These POVM elements span a subspace $\mathcal{V}_{D}=\mathrm{Span}\{E_{k}\}_{k=1}^{n}$ in the Hilbert-Schmidt space (HS) of the dimension $|\mathcal{V}_{D}|=D$.
 Typically, $D=d^{2}$ (the entire HS space), but for the sake of generality, we shall keep $D\leq d^{2}$.
-Furthermore, we assume $D<n$, thus $E_{k}$ forms an over-complete basis in $\mathcal{V}_{D}$.
+Furthermore, we assume $D<n$, and thus $E_{k}$ forms an overcomplete basis in $\mathcal{V}_{D}$.
 The estimation task aims to approximate the mean values of an observable $A\in\mathcal{V}_{D}$, i.e., $\expectationvalue{A}=\mbox{Tr}\left[{A\rho}\right]$.
@@
 Similarly, the measured probabilities trivially satisfy $p_{k}=\mbox{Tr}\left[{E_{k}\rho}\right]=\mbox{Tr}\left[{E_{k}\bar{\rho}}\right]$.
-The post-processing task assigns a (matrix) estimator $\hat{\eta}\in\{\dots,\hat{\eta}_{k},\dots\}$, i.e.
-each outcome $E_{k}$ the corresponding matrix $\hat{\eta}_{k}$, which represents a single-shot approximation of $\bar{\rho}$.
-They can be calculated via an inversion map $\hat{\eta_{k}}=\mathcal{M}^{-1}(E_{k})$ Huang et al.
-(2020); Nguyen et al.
-(2022) which we will show is not unique, i.e.
-$\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$ with the latter (homogeneous) component containing auxiliary parameters that may be used as a free resource in post-processing.
-The overall classical shadow itself is obtained by averaging over the estimators obtained from $M$ random measurements: $M^{-1}\sum_{s}^{M}\hat{\eta}^{(s)}$.
-Assuming that all samples are i.i.d., the average value $\expectationvalue{\hat{\eta}}$ is immediately
+The postprocessing task assigns a (matrix) estimator $\hat{\eta}\in\{\dots,\hat{\eta}_{k},\dots\}$, i.e., for each outcome $E_{k}$, the corresponding matrix $\hat{\eta}_{k}$, which represents a single-shot approximation of $\bar{\rho}$.
+They can be calculated via an inversion map $\hat{\eta}_{k}=\mathcal{M}^{-1}(E_{k})$ [1, 27], which we will show is not unique, i.e., $\mathcal{M}^{-1}=\mathcal{M}_{p}^{-1}+\mathcal{M}_{h}^{-1}$, with the latter (homogeneous) component containing auxiliary parameters that may be used as a free resource in postprocessing.
+The overall classical shadow itself is obtained by averaging over the estimators obtained from $K$ random measurements: $K^{-1}\sum_{s=1}^{K}\hat{\eta}^{(s)}$.
+Assuming that all samples are independent and identically distributed, the average value $\expectationvalue{\,\hat{\eta}\,}$ is immediately
 $\expectationvalue{\,\hat{\eta}\,}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}=\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
@@
 With Eq.
-(3), the task of finding an inversion map can be made equivalent to identifying a dual frame to the POVM in space of operators Innocenti et al.
-(2023).
-Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, the solution to Eq.
+(3), the task of finding an inversion map can be made equivalent to identifying a dual frame to the POVM in space of operators [40].
+Since the POVM forms an overcomplete basis in $\mathcal{V}_{D}$, the solution to Eq.
 (3) is not unique.
-This multiplicity results in a set of free parameters, which becomes an extra resource that may be used in the post-processing component of an estimation to reduce the variance.
+This multiplicity results in a set of free parameters, which becomes an extra resource that may be used in the postprocessing component of an estimation to reduce the variance.
 To achieve this, we now show a constructive method to recover families of equivalent classical shadows in terms of the POVM elements and their free parameters.
 We rely on an equivalent problem to Eq.
-(3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
+(3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$,
 $\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]=\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
@@
 where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$ is the inner product in $\mathcal{V}_{D}$.
-To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$
+To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$,
 $\sum_{k=1}^{n}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|=\sum_{k=1}^{n}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|\equiv\mathbbm{1}_{D},$ (6)
@@
 with $L_{sk}=\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle$ and $R_{tk}=\langle\!\langle{B_{t}}|{E_{k}}\rangle\!\rangle$.
-This implies the matrix equation $LR^{\dagger}=\mathbbm{1}_{D}$; thus, $L$ is a generalised inverse of $R^{\dagger}$.
-Since $R$ is a $D\times n$ matrix having $|{E_{k}}\rangle\!\rangle$ as columns, it is a rank-$D$ matrix (recall there are only $D$ independent $E_{k}$ matrices).
-We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero-matrix.
-Having this, a simple exercise shows $L=U(\Sigma^{-1}|H)W^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix.
-Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=U(\Sigma^{-1}|\hat{0})W^{\dagger}$ (particular) and $L^{(h)}=U(\hat{0}|H)W^{\dagger}$ (homogeneous).
-This results in a family of equivalent estimators
+This implies the matrix equation $LR^{\dagger}=\mathbbm{1}_{D}$; thus, $L$ is a generalized inverse of $R^{\dagger}$.
+Since $R$ is a $D\times n$ matrix having $|{E_{k}}\rangle\!\rangle$ as columns, it is a rank-$D$ matrix (recall there is only a number $D$ of independent $E_{k}$ matrices).
+We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero matrix.
+Having this, a simple exercise shows $L=U(\Sigma^{-1}|H)W^{\dagger}$, with $H$ being an arbitrary $D\times(n-D)$ matrix.
+Accordingly, we can split the solution into two parts, $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=U(\Sigma^{-1}|\hat{0})W^{\dagger}$ (particular) and $L^{(h)}=U(\hat{0}|H)W^{\dagger}$ (homogeneous).
+This results in a family of equivalent estimators,
 $\hat{\eta}_{k}=\sum_{s=1}^{D}L_{sk}\hat{B}_{s}=\sum_{s=1}^{D}\left(L_{sk}^{(p)}+L_{sk}^{(h)}\right)\hat{B}_{s}=\hat{\eta}_{k}^{(p)}+\hat{\eta}_{k}^{(h)},$ (8)
-parameterised by a homogeneous solution $\hat{\eta}_{k}^{(h)}$.
-As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimisation in post-processing.
+parameterized by a homogeneous solution $\hat{\eta}_{k}^{(h)}$.
+As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimization in postprocessing.
 We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$.
 The homogeneous coefficients, from the definition of $\hat{\eta}_{k}^{(h)}$ in Eq.
-(8), do not influence expectation values or any other linear combination of classical shadows
+(8), do not influence the expectation values or any other linear combination of classical shadows
 $\expectationvalue{\,a\,}=\sum_{k=1}^{n}p_{k}a_{k}=\sum_{k=1}^{n}p_{k}\left(a_{k}^{(p)}+a_{k}^{(h)}\right)=\sum_{k=1}^{n}p_{k}a_{k}^{(p)}.$ (9)
@@
 $\sum_{k}\,a_{k}^{(h)}\,E_{k}=0.$ (11)
-The set of solutions to the latter (homogeneous) equation defines an auxiliary optimisation space of dimension $n-D$ independent of the target observable (fixed by the choice of POVM).
-This is a free resource that may now be used in post-processing.
+The set of solutions to the latter (homogeneous) equation defines an auxiliary optimization space of dimension $n-D$ independent of the target observable (fixed by the choice of POVM).
+This is a free resource that may now be used in postprocessing.
 Figure 2: (a) Optimal free parameter for projectors on the Bloch sphere.
-The contour plot shows the minimal shadow norm Eq.
-(13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$).
-Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map Eq.
-(15) found in Huang et al.
-(2020) (b) Optimal free parameter for plane projectors.
+The contour plot shows the minimal shadow norm, given by Eq.
+(13), of the single-qubit projector characterized by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$).
+Optimization results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map, given by Eq.
+(15), found in [1].
+(b) Optimal free parameter for plane projectors.
 The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$).
@@
 In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis, the bound achieves optimal value |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
-Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy.
-It is customary in classical shadow tomography to refer to variance as figure of merit Huang et al.
-(2020).
-In particular, we consider the state-independent upper bound on variance, called shadow norm $\|A\|^{2}_{\text{sh}}$, defined as:
+Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is, the expected sample size of copies required to estimate a property of the state up to a fixed accuracy.
+It is customary in classical shadow tomography to refer to variance as the figure of merit [1].
+In particular, we consider the state-independent upper bound on variance, called shadow norm $\|A\|^{2}_{\text{sh}}$, defined as
 $\begin{split}\mbox{Var}\left[{a}\right]&=\mathbb{E}[a^{2}]-(\mathbb{E}[a])^{2}\leq\mathbb{E}[a^{2}]=\sum_{k=1}^{n}p_{k}|a_{k}|^{2}\\ &\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{sh}}.\end{split}$ (12)
-Considering the upper bound to study sample complexity is reasonable since, for bounded observables, the quantity $(\mathbb{E}[a])^{2}$ is also bounded, therefore the relevant contribution is brought by the first term $\mathbb{E}[a^{2}]$.
+Considering the upper bound to study sample complexity is reasonable since, for bounded observables, the quantity $(\mathbb{E}[a])^{2}$ is also bounded, and therefore the relevant contribution is brought by the first term $\mathbb{E}[a^{2}]$.
 This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably.
-This is akin to the original upper bound on shadow norm and thence sampling complexity found in Huang et al.
-(2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{sh}})$ ).
-Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator.
-By optimising over these coefficients, it is possible to identify the configuration which guarantees the minimal sample size for the estimation of the particular observable $A$:
+This is akin to the original upper bound on the shadow norm and thence the sampling complexity found in [1], where the expected number of samples scales with $M=O(\varepsilon^{-2}\ln\,{\delta^{-1}}\|A\|^{2}_{\text{sh}})$.
+Nevertheless, this can be further optimized over the free parameters in the homogeneous part of the estimator.
+By optimizing over these coefficients, it is possible to identify the configuration which guarantees the minimal sample size for the estimation of the particular observable $A$,
 $\|A\|^{2}_{\text{opt}}=\min_{a_{k}^{(h)}:H}\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right],$ (13)
 The existence of a minimum is guaranteed through the positive semidefiniteness of the variance operator.
-We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in Eq.
+We will provide examples to illustrate that different choices of the free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimization in Eq.
 (13) yields exceptional gains for relatively minuscule cost.
-Product observables and shadow optimisation
+Product observables and shadow optimization
 Let us now discuss the estimation of product observables $A^{(N)}=A_{1}\otimes A_{2}\ldots\otimes A_{N}$ via local POVMs, i.e., of the type $E_{k_{1}}\otimes\dots\otimes E_{k_{n}}$.
-This case is particularly interesting for near-term applications, such as the VQE algorithm Peruzzo et al.
-(2014); Tilly et al.
-(2022).
+This case is particularly interesting for near-term applications, such as the Variational Quantum Eigensolver (VQE) algorithm [42, 43].
 In such a scenario, it is easy to show that the shadow norm in Eq.
-(12) has a product form (see also Nguyen et al.
-(2022)) and consequently Eq.
+(12) has a product form (see, also, [27]) and, consequently, Eq.
 (13) becomes
 $\|A^{(N)}\|^{2}_{\text{opt}}=\prod_{j=1}^{N}\|A_{j}\|^{2}_{\text{opt}}$ (14)
-Since we are considering local measurements on independent subsystems, each term $\|A_{j}\|^{2}_{\text{opt}}$ can be optimised locally, implying the optimisation in Eq.
+Since we are considering local measurements on independent subsystems, each term $\|A_{j}\|^{2}_{\text{opt}}$ can be optimized locally, implying the optimization in Eq.
 (14) is efficient overall.
 We see that even minimal deviation from the minimum in local shadow norms $\|A_{j}\|^{2}_{\text{opt}}$ may accumulate exponentially fast with $N$.
-This is illustrated in Figure 1(b) for the case of the canonical inversion map of Eq.
+This is illustrated in Fig.
+1(b) for the case of the canonical inversion map of Eq.
 (15).
@@
 We first consider the case of estimation of fidelity with product states.
-The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterised on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$.
-The measurement scheme is described by the POVM of normalised projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{6}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$.
+The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterized on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$.
+The measurement scheme is described by the POVM of normalized projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{6}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$.
 Using the estimators
 $\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}\right).$ (15)
-found in Huang et al. (2020), the shadow norm in Eq. (12) simply evaluates for the generic single-qubit observable $A=a\cdot\mathbbm{1}+x\cdot\sigma_{x}+y\cdot\sigma_{y}+z\cdot\sigma_{z}$ to
+found in [1], the shadow norm in Eq. (12) simply evaluates for the generic single-qubit observable $A=a\cdot\mathbbm{1}+x\cdot\sigma_{x}+y\cdot\sigma_{y}+z\cdot\sigma_{z}$ to
 $\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (16)
-Considering the particular case of projectors on the Bloch sphere, with the parameterisation
+Considering the particular case of projectors on the Bloch sphere, with the parameterization
 $\begin{split}a\to\frac{1}{2},\quad\quad x\to\frac{1}{2}\sin\theta\cos\phi,\\ y\to\frac{1}{2}\sin\theta\sin\phi,\quad\quad z\to\frac{1}{2}\cos\theta,\end{split}$ (17)
@@
 Now we show how to improve this reference bound by our techniques.
-The POVMs $E^{\pm}_{\zeta}$ span the entire HS space and we may determine easily a homogeneous combination of the effects as in Eq.
+The POVMs $E^{\pm}_{\zeta}$ span the entire HS space and we may easily determine a homogeneous combination of the effects as in Eq.
 (11) given $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$.
@@
 As expected, we find $n-D=2$ parameters $p_{1},p_{2}$.
-Optimising over these parameters, we present the minimal upper bound as a function of angles $(\theta,\phi)$ in Fig.2(a).
-The optimised results are contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$.
-Since the overall shadow norm Eq.
-(14) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
+Optimizing over these parameters, we present the minimal upper bound as a function of angles $(\theta,\phi)$ in Fig .2 (a).
+The optimized results are contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$.
+Since the overall shadow norm given by Eq.
+(14) is of the product form, there are entire regions in the $(\theta,\phi)$ space for which deviations from the standard $3/2$ lead to the exponential improvement in $N$.
 Estimation with constrained POVMs
@@
 with $D=d^{2}$.
-Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an over-complete basis (within $\mathcal{V}_{D}$).
-Note that the standard inversion procedure Huang et al.
-(2020); Nguyen et al.
-(2022) is defined on the full HS space and does not directly apply in this instance.
-Consider for instance the POVM constituted by the projectors of the X and Y Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the ”equator” on the Bloch sphere.
-With a straightforward application of our procedure Eq.
+Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an overcomplete basis (within $\mathcal{V}_{D}$).
+Note that the standard inversion procedure [1, 27] is defined on the full HS space and does not directly apply in this instance.
+Consider, for instance, the POVM constituted by the projectors of the $X$ and $Y$ Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the “equator” on the Bloch sphere.
+With a straightforward application of our procedure given by Eq.
 (8), we obtain the following particular solution
@@
 The homogeneous term is added by solving $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$, where $\{h_{\zeta}\}_{\zeta=x,y}=\{p,-p\}$ and $p$ is a free parameter .
-For the generic planar observable (within span of $\{E_{\zeta}^{\pm}\}$)
+For the generic planar observable (within the span of $\{E_{\zeta}^{\pm}\}$),
 $A=\matrixquantity(a&x+iy\\ x-iy&a),$ (20)
-we find the shadow norm
+we find the shadow norm,
 $\|A\|^{2}_{\text{sh}}(p)$ $=(a-x)^{2}+2(x^{2}+y^{2})+B(p)+\sqrt{C(p)},$ (21)
@@
 […]
-We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies Huang et al. (2020).
+We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies [1].
 Optimal estimation of planar Paulis
-When considering as target observable the planar Pauli,
+When considering, as the target observable, the planar Pauli,
 $a\to 0,\quad x\to\cos\varphi,\quad y\to\sin\varphi,$ (22)
-the optimised shadow norm from Eq.
+the optimized shadow norm from Eq.
 (21) results in $\|A\|^{2}_{\text{opt}}=2$.
-An immediate result is that the expected number of resources for product of planar Paulis scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements Huang et al.
-(2020), $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$.
-Surprisingly, we are able to match the optimal scaling Huang et al.
-(2021b) while still relying only on local single-qubit measurements.
+An immediate result is that the expected number of resources for the product of planar Pauli eigenstates scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements [1], $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$.
+Surprisingly, we are able to match the optimal scaling [44] while still relying only on local single-qubit measurements.
 Complete versus overcomplete POVMs
 In the context of the previous example, it is interesting to compare the performance of complete versus overcomplete POVM.
-For this purpose, consider the example of a triangle POVM on the plane, defined as the normalised projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
+For this purpose, consider the example of a triangle POVM on the plane, defined as the normalized projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
 $|{\psi_{k}}\rangle=\dfrac{|{0}\rangle+e^{i\frac{k}{3}2\pi}|{1}\rangle}{\sqrt{2}}\quad k=0,1,2$ (23)
@@
 Since the POVM is complete, there is no free parameter dependence.
-From here, we can compute the shadow norm as defined in Eq.(12).
-For a planar Pauli as target observable, as in Eq.
+From here, we can compute the shadow norm as defined in Eq.
+(12).
+For a planar Pauli as the target observable, as in Eq.
 (22), the shadow norm is $\|A\|^{2}_{\text{sh}}=3,\,\forall\varphi$, which leads to an expected overhead of $3^{N}$ on the scaling of resources.
 From the previous example, the same observables can be optimally estimated with an overhead of $2^{N}$.
-We have therefore found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
+We have, therefore, found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
 Locally efficient fidelity estimation implies globally efficient
-We now present a configuration in which the optimisation of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e.
-$\|A\|^{2}_{\text{opt}}=1$.
-To show this, we choose as target observables tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ | where each single-qubit state is parameterised as: $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
-Using the change of coordinates
+We now present a configuration in which the optimization of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e., $\|A\|^{2}_{\text{opt}}=1$.
+To show this, we choose, as target observables, tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ |, where each single-qubit state is parameterized as $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
+Using the change of coordinates,
 $a\to\frac{1}{2},\quad x\to\frac{1}{2}\cos\phi,\quad y\to\frac{1}{2}\sin\phi,$ (25)
-we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$:
+we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$,
 |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}=\frac{1}{2}$ $(1+p_{i}^{2}+B(\phi,p_{r})+$ (26)
@@
 This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $4^{N}$ projections simultaneously for complementary observables.
-Previously this result has been found in Huang et al.
-(2020) only via global measurement strategies relying on $N-$qubit Clifford unitaries.
+Previously, this result has been found in [1] only via global measurement strategies relying on $N-$qubit Clifford unitaries.
 We have shown instead that this is achieved using exclusively local measurements.
 Discussion
-In the course of this letter we have observed how post-processing, an often overlooked aspect in estimation, represents a relevant resource to be examined in the accounting of a more efficient tomography.
-By introducing a generalised framework to obtain families of equivalent parameter-dependent estimators, we have found a clear relation between over-complete POVMs and their ability to reduce the variance in estimation of arbitrary observables.
-When considering local measurements on large composite quantum systems, even a small gain on each estimator represents an overall improvement in sample complexity, yielding an exponential separation between the naive sample variance and the optimised one as seen in Figure 1(b).
-In the case of example two, this optimised gain can achieve a variance that performs as well as global classical shadows for given observables, with modest additional post processing that remains efficient.
-We anticipate that by considering overcomplete POVMs containing joint measurements on multiple qubits, further advantage may be found at increased (but still tractable) optimisation difficulty, either in sample complexity or the set of observables.
-Finally, interleaving optimised multiqubit measurements may yield sample improvements that approach regular classical shadows within some limit.
-Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimisation will be the topic of future work.
-Note - In preparation of this manuscript, we became aware of two other works Malmi et al.
-(2024); Fischer et al.
-(2024) considering a similar problem.
-The authors consider a similar optimisation of the post-processing stage in estimation tasks, albeit referring to different frameworks.
+In the course of this paper, we have observed how postprocessing, an often overlooked aspect in estimation, represents a relevant resource to be examined in the accounting of a more efficient tomography.
+By introducing a generalized framework to obtain families of equivalent parameter-dependent estimators, we have found a clear relation between overcomplete POVMs and their ability to reduce the variance in the estimation of arbitrary observables.
+When considering local measurements on large composite quantum systems, even a small gain on each estimator represents an overall improvement in sample complexity, yielding an exponential separation between the naive sample variance and the optimized one as seen in Fig.
+1(b).
+In the case of example two, this optimized gain can achieve a variance that performs as well as global classical shadows for given observables, with modest additional post processing that remains efficient.
+We anticipate that by considering overcomplete POVMs containing joint measurements on multiple qubits, further advantage may be found at increased (but still tractable) optimization difficulty, either in sample complexity or the set of observables.
+Finally, interleaving optimized multiqubit measurements may yield sample improvements that approach regular classical shadows within some limit.
+Such questions are natural candidates for further investigation and the identification of such a hierarchy of optimization will be the topic of future work.
+Note added - Recently, we became aware of two other works [45, 46] considering a similar problem.
+The authors consider a similar optimization of the postprocessing stage in estimation tasks, albeit referring to different frameworks.
 Acknowledgements
 We thank Richard Küng for his insight and excellent discussions on the subject matter.
-This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/F71] (BeyondC).
-For the purpose of open access, the author(s) has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
+This research was funded in whole or in part by the Austrian Science Fund (FWF)(Grant No.10.55776/F71]) (BeyondC).